# Robust Hallucination Detection in LLMs via Adaptive Token Selection

### v1

@@ Abstract
Hallucinations in large language models (LLMs) pose significant safety concerns that impede their broader deployment. Recent research in hallucination detection has demonstrated that LLMs’ internal representations contain truthfulness hints, which can be harnessed for detector training. However, the performance of these detectors is heavily dependent on the internal representations of predetermined tokens, fluctuating considerably when working on free-form generations with varying lengths and sparse distributions of hallucinated entities. To address this, we propose HaMI, a novel approach that enables robust detection of hallucinations through adaptive selection and learning of critical tokens that are most indicative of hallucinations. We achieve this robustness by an innovative formulation of the Hallucination detection task as Multiple Instance (HaMI) learning over token-level representations within a sequence, thereby facilitating a joint optimisation of token selection and hallucination detection on generation sequences of diverse forms. Comprehensive experimental results on four hallucination benchmarks show that HaMI significantly outperforms existing state-of-the-art approaches.
@@ 1 Introduction
Recent progress in Large Language Models (LLMs) has demonstrated impressive capabilities across a wide range of applications. However, the ever-growing popularity of LLMs also gives rise to concerns about the reliability of their outputs Ji et al. (2023); Chuang et al. (2023). Some research has indicated that LLMs are susceptible to hallucinations, which can be described as unfaithful or incorrect generations Longpre et al. (2021); Adlakha et al. (2023); Zhang et al. (2023). This tendency not only impedes the broader applications of LLMs but also poses potential safety risks, especially in high-stake fields such as legal and medical services. Therefore, the reliable detection of hallucinations is critical for the safe deployment of LLMs.
Various approaches have been developed to detect hallucinations. Recent studies indicate that predictive uncertainty can serve as useful detection features Kadavath et al. (2022); Burns et al. (2022); Duan et al. (2024), as predictions with low confidence often correlate with the presence of hallucinated content. Some research focuses on the evaluation of a single generation Manakul et al. (2023) while some harness semantic information from several samples, e.g., the semantic equivalences across multiple generations for the same question Farquhar et al. (2024). The latter is to exploit the observation that LLMs would consistently generate the same response if they were certain of it. Nevertheless, these methods are mainly based on the final generation output, rendering them ineffective in leveraging important semantics in the internal state representations.
@@
Another line of research focuses on the utilisation of the internal states of LLMs for hallucination detection. These internal representations can encode significant information about truthfulness direction of the generations. Motivated by this, great efforts have been made to explore the characteristics of these representations and train a binary classifier on them. Various supervision signals, including accuracy labels Li et al. (2024), converted semantic entropy labels Kossen et al. (2024), and eigenvalue-related labels Du et al. (2024), have been harnessed to train the classifier for detection tasks. One major challenge for these methods is that the majority of tokens in an incorrect/hallucinated response may not contribute to truthfulness. To address this issue, most of these methods focus on the use of predetermined tokens, such as the first generated token, the last generated token, or the one before the last. However, the exact location of the most indicative tokens for hallucination can vary significantly for generations of different questions, as illustrated in Figure 1, since the generation responses are of free-form with varying lengths and have sparse distributions of hallucinated entities. As a result, they can overlook important tokens where hallucinated information is actually concentrated.
To address this challenge, we propose HaMI, namely, Hallucination detection as Multiple Instance learning, an end-to-end joint token selection and hallucination detection approach that enables adaptive token selection from internal state representations for stable and accurate hallucination detection on generation responses of varying length. In HaMI we reformulate the task as a multiple instance learning (MIL) problem Carbonneau et al. (2018), where each response sequence is treated as a bag of token instances, with a bag-level label as either hallucinated (positive) or trustworthy (negative), and the objective becomes binary classification of the token bags. This way takes advantage of the fact that only a few token instances in the positive bag are positive since hallucinated content typically manifests in only a small subset of tokens within a sequence, whereas all token instances in the negative bag are always negative. In doing so, the MIL approach enables the exploitation of the hallucination labels at the sequence (bag) level to adaptively select the most responsive tokens for sequence-level hallucination detection.
To be more specific, LLMs are first prompted to generate response sequences with varying length. The MIL-driven hallucination detector is then optimised in HaMI to assign hallucination scores to all individual token instances and adaptively select the most indicative tokens in both positive and negative bags for the sequence-level prediction. The optimisation results in a detector that can distinguish the selected most positive hallucinated token instances from the hard negative instances (i.e., the tokens that have highest hallucination scores in a negative bag). Additionally, recognizing that predictive uncertainty serves as an important indicator of correctness, we further propose a representation enhancement module in HaMI, where we integrate multiple levels of uncertainty information into the original representation space for more effective training of our HaMI detector.
@@
  • We propose a novel MIL-based framework HaMI for hallucination detection, which enables an end-to-end joint optimisation of token selection and hallucination detection. This results in adaptive hallucinated token selection, effectively mitigating the performance instability on response generations of varying length and hallucination entities. To our best knowledge, this is the first approach allowing such a joint optimisation.
  • We further introduce a module that incorporates internal representations with uncertainty scores to provide more hallucination indication information for the join optimisation in HaMI.
  • Comprehensive empirical results with widely adopted LLMs on four popular benchmark datasets show that HaMI can significantly outperform state-of-the-art (SOTA) competing methods.
Figure 2: The framework of our proposed HaMI. The LLM is prompted to generate answer tokens accompanied by token representations $h_{i}$. The network receives sequences of token representations from both positive $\mathcal{B^{+}}$ and negative $\mathcal{B^{-}}$ bags as inputs. The hallucination detector initially extracts high-level representations, followed by assigning a hallucination score to each token instance. We choose the largest scores from both positive and negative bags, subsequently maximizing the discriminative margin between them by minimizing a MIL loss as described in Eq. 4. Given the sequential nature of the generations, a constraint on the smoothness of hallucination scores of adjacent tokens is also added into HaMI.
@@ 2 Related Work
The term hallucination, under the closed-book setting, can refer to unfaithful or fabricated generations Zhang et al. (2023); Ji et al. (2023). Having gained wide interest, hallucination detection is crucial for LLMs to maintain high reliability in various specific tasks. These methods can be categorised as two main lines: uncertainty measurement and internal state analysis.
Uncertainty measurement.
Uncertainty measurement has been widely explored for hallucination detection. Some research focuses on token-level uncertainty Talman et al. (2023); Duan et al. (2024) with the assumption that low predictive logits or high entropy over tokens’ predictive distribution indicates the high possibility of hallucination. Some studies quantify sentence-level uncertainty by instructing LLMs to express the predictive uncertainty themselves with prompts like “Is your answer True or False?” Kadavath et al. (2022); Lin et al. (2022); Zhou et al. (2023); Manakul et al. (2023). Furthermore, there are many studies exploring exploiting semantic equivalence by measuring the consistency among multiple responses sampled from LLMs Mündler et al. (2023); Dhuliawala et al. (2023). For example, Farquhar et al. (2024) proposes Semantic Entropy, which employs a powerful LLM to evaluate semantic entailment among multiple generations and calculate semantic entropy over the entailment identity as uncertainty scores. Nonetheless, the performance of these methods relies on external tools and ignores the important semantics embedded in the internal representations of LLMs.
Internal state analysis.
@@
Recently, a branch of works suggests that the internal states of LLMs encode more knowledge than they express and can reveal truthfulness direction Hubinger et al. (2024); Chen et al. (2024). The majority of this line employs probes Alain (2016) to better understand layer-wise representations and predict the correctness of generations Li et al. (2024); Marks and Tegmark (2024). Some research extends these methods by proposing new supervision signals. For example, Du et al. (2024) performs singular value decomposition on internal representations and calculates the norm of these representations projected on singular vectors as class scores, which are converted to class labels based on insights from a small set of wild data. Kossen et al. (2024) argue that the aforementioned semantic entropy value is preferable to accuracy labels for supervised training. While most of these works leverage predefined token representations, the truthfulness information is concentrated in specific tokens. Orgad et al. (2024) attempt to prompt an LLM to find the exact token in the sequence, but this method is resource-intensive and the reliability of detection is greatly impressed by the capability of the employed LLM. Instead of employing another LLM to find the unique exact answer, we train a ranking model to automatically select the critical tokens for hallucination detection in an end-to-end manner.
@@ 3 Preliminaries
Given a sequence of input tokens $\mathbf{x}=\{x_{1},x_{2},\dots,x_{n}\}$ consisting of a specific question $q$, an LLM will generate a sequence of tokens $\mathbf{y}=\{y_{1},y_{2},\dots,y_{t}\}$. Generally, each token $y_{i\in\{1,2,\dots,t\}}$ is decoded from the next predictive distribution over the model’s vocabulary set $\mathcal{V}$, formulated as $y_{i}=\arg\max_{y\in\mathcal{V}}P(y|y_{<i},\mathbf{x})$, and the predictive probability, i.e., logit, is denoted by $p_{i}$ for short. For each $\mathbf{y}$, apart from the token level uncertainty $p_{i}$, we can also obtain the overall uncertainty of the entire sequence, presented as
$p_{s}=(\prod_{i=1}^{t}p_{i})^{-\frac{1}{t}}.$ (1)
By accessing the internal states of the model, we can extract the internal representation $h_{i,l}\in\mathbbm{R}^{d}$ at layer $l$ for each token $y_{i}$, where $d$ is decided by the dimensions of the internal states of the LLM. The sequence feature space is denoted as $\mathcal{H}.$ The correctness of generated response is evaluated by GPT-4 Achiam et al. (2023) with label $z\in\{0,1\}$.
Given a datset $D=\{(q_{n},a_{n}\}_{n=1}^{N}$, where $\{q_{n}\}_{n=1}^{N}$ and $\{a_{n}\}_{n=1}^{N}$ are questions and ground-truth answers respectively, LLMs will generate answers $\mathbf{y}_{n}$ accompanied with predictive logits, hidden representations =$\mathcal{H}=\{\mathbf{h}_{n}\}_{n=1}^{N}$ and correctness labels $z_{n}\in\{0,1\}$ evaluated by GPT-4 Achiam et al. (2023). Our MIL-based hallucination detection is to automatically identify the most representative tokens in incorrect responses and the hard negative tokens (i.e., the most likely hallucinated tokens within correct responses) for training stable and accurate detectors.
@@ 4 Methodology
Our proposed HaMI aims to distinguish hallucination-free and hallucination-containing text generations by an adaptive token selection approach implemented by a joint optimisation of token selection and hallucination detection. Additionally, we introduce a predictive uncertainty-based module to integrate more hallucination features in HaMI and enhance its discriminative capability. The overall framework is presented in Figure 2. Below we introduce these two modules in detail.
@@ 4.1 MIL-driven Adaptive Token Selection
For a given sentence, its correctness is often determined by just a few words, such as noun entities Chuang et al. (2023), suggesting that truthfulness information may be encoded in the internal representations of specific tokens. Nevertheless, the correctness label is typically assigned to the entire sequence. The key idea of our approach is to adaptively identify the most salient tokens, upon which we can train a reliable hallucination detector. To achieve this goal, we introduce Multiple Instance Learning (MIL) to this domain.
@@
Therefore, we reformulate hallucination detection with adaptive token selection as a MIL problem. Particularly, the generated sequence can be regarded as a bag of tokens. Generations without hallucinations are labelled as negative bags $\mathcal{B^{-}}$ (label ‘0’), while those with hallucinations are labelled as positive bags $\mathcal{B^{+}}$ (label ‘1’). The representations of the positive and negative sequences are denoted as $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$ respectively and $\mathbf{h}=\{h_{i}\}_{i=1}^{t}$. Intuitively, the detector is expected to assign higher scores to token instances from the positive bag compared to those from the negative bag. Given the hallucination sparsity insight mentioned before, we choose the token instance with the highest score as the salient token, which can be defined as
$i^{+}=\arg\max\limits_{i}f_{\theta}(\mathbf{h}^{+}),$ (2)
$i^{-}=\arg\max\limits_{i}f_{\theta}(\mathbf{h}^{-}),$ (3)
where $i\in\{1,2,\cdots,t\}$, $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$ are drawn from input representation space $\mathcal{H}$ w.r.t. $\mathcal{B^{-}}$ and $\mathcal{B^{-}}$ respectively. As such, we are able to locate the instance most likely to represent a true positive in the positive bag and the most challenging instance resembling a hallucination in the negative bag as aforementioned salient tokens. Our approach seeks to distinguish between positive and negative samples by maximizing the distance between the selected instances from these two categories in representation space. The training MIL objective can then be formulated in a hinge-loss style as follows:
$\mathcal{L}_{MIL}=max(0,1-\max_{n\in\mathcal{B^{+}}}f_{\theta}(\mathbf{h}_{n}^{+})+\max_{n\in\mathcal{B^{-}}}f_{\theta}(\mathbf{h}_{n}^{-}),$ (4)
where $\theta$ are parameters of the hallucination detector as presented in Figure 2. In doing so, the adaptive token selection can mitigate the limitations caused by predefined token location. For example, the commonly used First generated token is hard to capture hallucination appears in the subsequent generations, especially when the generation is long or LLMs react to the input politely.
Note that the next token is conditioned on all previous tokens, so the token generation process contains a sequential nature. As depicted in Figure 2, on the same side of the peak, the hallucination scores for two adjacent tokens tend to be more similar. Therefore, we exploit sequential smoothness via the following loss
@@
through which we aim to ensure the consistency of the hallucination scores of neighbouring tokens. Therefore, the adaptive token selection is achieved by minimizing the following overall loss:

### v2

@@ Abstract
Hallucinations in large language models (LLMs) pose significant safety concerns that impede their broader deployment. Recent research in hallucination detection has demonstrated that LLMs’ internal representations contain truthfulness hints, which can be harnessed for detector training. However, the performance of these detectors is heavily dependent on the internal representations of predetermined tokens, fluctuating considerably when working on free-form generations with varying lengths and sparse distributions of hallucinated entities. To address this, we propose HaMI, a novel approach that enables robust detection of hallucinations through adaptive selection and learning of critical tokens that are most indicative of hallucinations. We achieve this robustness by an innovative formulation of the Hallucination detection task as Multiple Instance (HaMI) learning over token-level representations within a sequence, thereby facilitating a joint optimisation of token selection and hallucination detection on generation sequences of diverse forms. Comprehensive experimental results on four hallucination benchmarks show that HaMI significantly outperforms existing state-of-the-art approaches. Code is available at https://github.com/mala-lab/HaMI.
@@ 1 Introduction
Recent progress in Large Language Models (LLMs) has demonstrated impressive capabilities across a wide range of applications. However, the ever-growing popularity of LLMs also gives rise to concerns about the reliability of their outputs ji2023survey ; chuang2023dola . Some research has indicated that LLMs are susceptible to hallucinations, which can be described as unfaithful or incorrect generations longpre2021entity ; adlakha2023evaluating ; huang2025survey . This tendency not only impedes the broader applications of LLMs but also poses potential safety risks, especially in high-stake fields such as legal and medical services. Therefore, the reliable detection of hallucinations is critical for the safe deployment of LLMs.
Various approaches have been developed to detect hallucinations. Recent studies indicate that predictive uncertainty can serve as useful detection features burns2022discovering ; kadavath2022language ; duan-etal-2024-shifting , as predictions with low confidence often correlate with the presence of hallucinated content. Some research focuses on the evaluation of a single generation manakul2023selfcheckgpt ; bakman-etal-2024-mars while some harness the information contained in multiple generations. Compared to single-generation methods, approaches utilising multiple generations demonstrate greater effectiveness since they capture a broader set of hallucination signals. Such signals include semantic consistency farquhar2024detecting for the repeated queries, as well as behaviour inconsistency when LLMs are presented with a set of unrelated follow-up questions pacchiardi2024how after the initial question. Nevertheless, these methods are mainly based on the final generation output, rendering them ineffective in leveraging important semantics in the internal representations.
In parallel, another line of research focuses on the utilisation of the internal representations of LLMs for hallucination detection azaria-mitchell-2023-internal ; liu2025attention . These internal representations can encode information about the truthfulness direction of the generations. Motivated by this, great efforts have been made to explore the characteristics of these representations, most commonly by training a binary classifier on them. Various supervision signals, including accuracy labels li2024inference , converted semantic entropy labels kossen2024semantic , and eigenvalue-related labels du2024haloscope , have been harnessed to train the classifier for detection tasks. One major challenge for these methods is that a majority of tokens in an incorrect/hallucinated response may not contribute to truthfulness. To address this issue, most methods utilise predetermined tokens, such as the first generated token, the last generated token, or the one before the last ji-etal-2024-llm ; barbero2025llms ; zucchet2025language . However, the exact location of the most indicative tokens for hallucination can vary significantly, as illustrated in Figure 1, since the generation responses are of free-form with varying lengths and have sparse distributions of hallucinated entities. As a result, they can overlook important tokens where hallucinated information is actually concentrated.
@@
To address this challenge, we propose HaMI, namely, Hallucination detection as Multiple Instance learning, a novel approach that jointly optimises token selection and hallucination detection in an end-to-end fashion. The joint optimisation enables adaptive token selection from internal state representations for stable and accurate hallucination detection on generation responses of varying length. In HaMI we reformulate the task as a multiple instance learning (MIL) problem carbonneau2018multiple , where each response sequence is treated as a bag of token instances, with a bag-level label as either hallucinated (positive) or trustworthy (negative), and the objective becomes binary classification of the token bags. This way takes advantage of the fact that only a few token instances in the positive bag are positive since hallucinated content typically manifests in only a small subset of tokens within a sequence, whereas all token instances in the negative bag are always negative. In doing so, the MIL approach enables the exploitation of the hallucination labels at the sequence (bag) level to adaptively select the most responsive tokens for sequence-level hallucination detection.
To be more specific, LLMs are first prompted to generate response sequences with varying length. The MIL-driven hallucination detector is then optimised in HaMI to assign hallucination scores to all individual token instances and adaptively select the most indicative tokens in both positive and negative bags for the sequence-level prediction. The optimisation results in a detector that can distinguish the most positive hallucinated token instances from the hard negative token instances (i.e.,, the tokens that have highest hallucination scores in a negative bag). Additionally, recognizing that predictive uncertainty serves as an important indicator of correctness, we further propose a representation enhancement module in HaMI, where we integrate multiple levels of uncertainty information into the original representation space for more effective training of our HaMI detector.
@@
  • We propose HaMI, a novel MIL-based framework for hallucination detection, which enables an end-to-end joint optimisation of token selection and hallucination detection. This warrants adaptive selection of hallucinated tokens that are optimal for the learned detector, effectively mitigating the performance instability on response generations of varying length. To our best knowledge, this is the first approach allowing such a joint optimisation.
  • We further introduce a module that incorporates internal representations with uncertainty scores to provide more indications of hallucination for the joint optimisation in HaMI.
  • Comprehensive empirical results with widely adopted LLMs on four popular benchmark datasets show that HaMI can significantly outperform state-of-the-art (SotA) methods.
@@ 2 Related Work
The term hallucination, under the closed-book setting, can refer to unfaithful or fabricated generations ji2023survey ; zhang2023siren . Having gained wide interest, hallucination detection is crucial for LLMs to maintain high reliability in various specific tasks. These methods can be categorised into two main lines: uncertainty measurement and internal representation analysis.
Uncertainty Measurement-based Methods.
Uncertainty measurement has been widely explored for hallucination detection. Some research focuses on token-level uncertainty talman-etal-2023-uncertainty ; duan-etal-2024-shifting ; quevedo2024detecting with the assumption that low predictive logits or high entropy over tokens’ predictive distribution indicate a high possibility of hallucination. Some studies target sentence-level uncertainty to capture a more holistic view of the whole response. Such uncertainty can be measured by (weighted) aggregation of all token logits bakman-etal-2024-mars ; ren2023outofdistribution or obtained by directly instructing LLMs to express the truthfulness of their generations with simple but effective prompts kadavath2022language ; lin2022teaching ; zhou2023navigating ; manakul2023selfcheckgpt . To further enrich semantic understanding and improve the detection accuracy, researchers also explore LLMs’ behaviour among multiple generations mundler2023self ; dhuliawala2023chain . For example, Farquhar et al. farquhar2024detecting propose Semantic Entropy as an evaluation of semantic consistency among multiple generations for hallucination detection, while Pacchiardi et al. pacchiardi2024how present that behaviour inconsistency over predefined follow-up unrelated questions can also indicate the correctness of the response to the initial question. Although effective, the performance of these methods relies on external tools and ignores the important semantics embedded in the internal representations of LLMs.
Internal Representation-based Methods.
@@
Recently, a branch of work suggests that the internal representations of LLMs encode more knowledge than they express and can reveal truthfulness direction yuksekgonul2023attention ; hubinger2024sleeper ; snyder2024early ; bayat2025steering . The majority of this line employs probes alain2016understanding to better understand head-wise or layer-wise representations and predict the correctness of generations burns2022discovering ; li2024inference ; chen2024inside ; marks2024geometrytruthemergentlinear . Recent research extends these methods by proposing new supervision signals, such as an automated membership estimation score presented in HaloScope du2024haloscope and the aforementioned semantic entropy value kossen2024semantic , which has been proven to be preferable to accuracy labels for supervised training. Most of these works leverage predefined token representations, but the truthfulness information is concentrated in specific tokens. Some research attempts to prompt an LLM to find the “exact token” in the sequence orgad2024llmsknowshowintrinsic , but the reliability of detection is greatly influenced by the capability of the employed LLM. Unlike existing methods that separate token selection and hallucination detection into two stages and require external assistance, we propose to train a ranking model to automatically select the critical tokens for hallucination detection in an end-to-end manner.
@@ 3 Preliminaries
Given a sequence of input tokens $\mathbf{x}=\{x_{1},x_{2},\dots,x_{m}\}$ consisting of a specific question $q$, an LLM will generate a sequence of tokens $\mathbf{y}=\{y_{1},y_{2},\dots,y_{t}\}$. Generally, each token $y_{i\in\{1,2,\dots,t\}}$ is decoded from the next predictive distribution over the model’s vocabulary set $\mathcal{V}$, formulated as $y_{i}=\arg\max_{y\in\mathcal{V}}P(y|y_{<i},\mathbf{x})$, and the predictive probability is denoted by $p_{i}$ for short. By accessing the internal states of the model, we can extract the internal representation $h_{i,l}\in\mathbbm{R}^{d}$ at layer $l$ for each token $y_{i}$, where $d$ is decided by the dimensions of the internal states of the LLM. The correctness of generated response is evaluated by GPT-4.1 achiam2023gpt with label $z\in\{0,1\}$.
Given a datset $D=\{(q_{n},a_{n})\}_{n=1}^{N}$, where $\{q_{n}\}_{n=1}^{N}$ and $\{a_{n}\}_{n=1}^{N}$ are questions and ground-truth answers respectively, LLMs will generate answers decoded from a token list $\mathbf{y}_{n}$ accompanied with predictive probabilities, hidden representations $\mathbf{h}_{n}$ and correctness labels $z_{n}\in\{0,1\}$. The representation space for all samples is denoted as =$\mathcal{H}=\{\mathbf{h}_{n}\}_{n=1}^{N}$ Our MIL-based hallucination detection is to identify the most representative tokens in incorrect responses and the hard negative tokens (i.e., the most likely hallucinated tokens within correct responses) that are optimal for training the subsequent detector, i.e.,, a joint optimisation of token selection and hallucination detection.
Figure 2: The framework of our proposed HaMI. The LLM is prompted to generate answer tokens accompanied by token representations $h_{i}$. The network receives sequences of token representations for both positive $\mathcal{B^{+}}$ and negative $\mathcal{B^{-}}$ bags as inputs at the training stage. The hallucination detector assigns a hallucination score to each token instance. We choose the top $k$ largest scores from both bags, subsequently maximising the discriminative margin between them by minimising a MIL loss as described in Eq. 4. Given the sequential nature of the generations, a constraint on the smoothness of hallucination scores of adjacent tokens is also added to HaMI.
@@ 4 Methodology
Our proposed HaMI aims to distinguish hallucination-free and hallucination-containing text generations by adaptive token selection that jointly optimises token selection and hallucination detection. Additionally, we introduce a predictive uncertainty-based module to integrate more hallucination features in HaMI and enhance its discriminative capability. The overall framework is presented in Figure 2. Below we introduce these two modules in detail.
@@ 4.1 MIL-driven Adaptive Token Selection
For a given sentence, its correctness is often determined by just a few words, such as noun entities chuang2023dola , suggesting that truthfulness information may be encoded in the internal representations of specific tokens. Nevertheless, the correctness label is typically assigned to the entire sequence. The key idea of our approach is to adaptively identify the most salient tokens, upon which we can train a reliable hallucination detector. To this end, we introduce Multiple Instance Learning (MIL) to this domain.
@@
Therefore, we reformulate hallucination detection with adaptive token selection as a MIL problem. Particularly, the generated sequence can be regarded as a bag of tokens. Generations without hallucinations are labelled as negative bags $\mathcal{B^{-}}$ (label ‘0’), while those with hallucinations are labelled as positive bags $\mathcal{B^{+}}$ (label ‘1’). The representations of the positive and negative sequences are denoted as $\mathbf{h}^{+}$ and $\mathbf{h}^{-}$ respectively and $\mathbf{h}=\{h_{i}\}_{i=1}^{t}$. Intuitively, the detector is expected to assign higher scores to token instances from the positive bag compared to those from the negative bag. Given the hallucination sparsity insight mentioned before, we choose the token instances with the top $k$ highest hallucination scores as the salient tokens in each sequence, which can be defined as
$\mathcal{I}_{\text{top-}k}=\{i_{1},i_{2},\ldots,i_{k}\}\quad\text{s.t.}\quad\left\{\begin{array}[]{l}f_{\theta}(h_{i_{1}})\geq f_{\theta}(h_{i_{2}})\geq\cdots\geq f_{\theta}(h_{i_{k}})\\ f_{\theta}(h_{i_{k}})\geq f_{\theta}(h_{i_{j}}),\quad\forall i_{j}\notin\mathcal{I}_{\text{top-}k}\end{array}\right.$ (3)
where $i_{k}\in\{1,2,\cdots,t\}$, and $f_{\theta}$ is the hallucination detector we aim to learn with the parameters $\theta$, as presented in Figure 2. As such, we are able to locate the most hallucinated token instances in the positive bag and the top hard negative token instances resembling a hallucination in the negative bag as the salient tokens. Our approach then seeks to distinguish between positive and negative bags by maximizing the distance between the selected token instances from these two categories in the representation space. The training MIL objective can be formulated as follows:
$\mathcal{L}_{MIL}=1-\lVert\frac{1}{k}\sum_{i^{+}\in\mathcal{I}_{\text{top-}k}^{+}}f_{\theta}(h_{i^{+}})\rVert_{2}+\lVert\frac{1}{k}\sum_{i^{-}\in\mathcal{I}_{\text{top-}k}^{-}}f_{\theta}(h_{i^{-}})\rVert_{2},$ (4)
where $h_{i^{+}}$ is selected from $\mathbf{h}^{+}_{n}\in\mathcal{B}^{+}$ and $h_{i^{-}}$ is from $\mathbf{h}^{-}_{n}\in\mathcal{B}^{-}$, and $\mathcal{I}_{\text{top-}k}^{+}$ and $\mathcal{I}_{\text{top-}k}^{-}$ are index sets for tokens in sequences from $\mathcal{B^{+}}$ and $\mathcal{B^{-}}$ respectively. In doing so, the adaptive token selection can mitigate the limitations caused by predefined token location.
Note that the next token is conditioned on all previous tokens, so the token generation process contains a sequential nature. As depicted in Figure 2, on the same side of the peak, the hallucination scores for two adjacent tokens tend to be more similar. Therefore, we exploit sequential smoothness via the following loss:
@@
where $i\in\{2,\cdots,t\}$. Through this, we aim to ensure the consistency of the hallucination scores of neighbouring tokens. Therefore, HaMI is optimised by minimizing the following overall loss: